# TD3Net: A Temporal Densely Connected Multi-Dilated Convolutional Network for Lipreading

### v1

@@ 6 Conclusion
In this study, we proposed TD3Net to maintain a wide and dense receptive field without blind spots for word-level lipreading, which is crucial for recognizing individual words and serves as the basis for recognizing longer sequences. TD3Net, which combines nested dense skip connections and multi-dilated TC layers, can learn multi-temporal representations in almost all layers without losing information about the continuity of lip movements induced by blind spots. Experiments on two public benchmark datasets revealed that the proposed approach achieves performance comparable to existing state-of-the-art methods. Notably, under tightly controlled conditions, TD3Net outperformed other TCN-based backend architectures while maintaining lower computational cost and fewer parameters. Moreover, we visually observed that standard dilation in dense skip connections created blind spots that disrupted temporal continuity, while our multi-dilation strategy preserved it. This temporal modeling is particularly important for lipreading systems—a claim further supported by quantitative results. Although TD3Net demonstrated strong performance in word-level recognition, future research is needed to extend it to sentence-level lipreading.

### v4

@@ 6 Conclusion
In this study, we proposed TD3Net to maintain a wide and dense receptive field without blind spots for word-level lipreading, which is crucial for recognizing individual words and serves as the basis for recognizing longer sequences. TD3Net, which combines nested dense skip connections and multi-dilated TC layers, can learn multi-temporal representations in almost all layers without losing information about the continuity of lip movements induced by blind spots. The experimental results on two public benchmark datasets revealed that the proposed approach achieves performance comparable to existing state-of-the-art methods. In particular, experiments conducted under tightly controlled conditions demonstrated that the proposed method yielded superior performance compared with other backend architectures while maintaining lower computational cost and fewer parameters. Moreover, we visually observed that standard dilation in dense skip connections created blind spots that disrupt temporal continuity, while our multi-dilation strategy preserved it. This capability is achieved in TD3Net through the dense integration of TD2 blocks, which stabilizes training and enhances temporal modeling, ultimately leading to significant improvements in lipreading performance. Although the proposed method demonstrated excellent lipreading performance, this study focused solely on recognizing individual words; therefore, further research is needed to develop a method that can accurately recognize speech at the sentence level in lipreading.